OpenStudy (anonymous):

what is square root -108 + swuareroot -27 ????

OpenStudy (anonymous):

use the equation generator please

OpenStudy (anonymous):

where is the equation generator

OpenStudy (anonymous):

the button that says equation on the other side of the post button

OpenStudy (anonymous):

Or you could just post it to wolframalpha since you're not really trying to get help as much as just get the answer.

OpenStudy (anonymous):

\[ \sqrt{-6}(2+ \sqrt{-8})\]

OpenStudy (anonymous):

can you explain how to do it if yu can please and try to ignore polpak i am trying to get help hence on y i am posting questions like everyone else

OpenStudy (anonymous):

you know that \[\sqrt{i }=\sqrt{-1}\] and that \[\sqrt{ab}=\sqrt{a}\sqrt{b}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiply the \[\sqrt{-6}\] and show me what you get

OpenStudy (anonymous):

so your saying break down the -6 and then muliply it by the whole equation

OpenStudy (anonymous):

multiply the \[\sqrt{-6} \] in first then show me what you get

OpenStudy (anonymous):

\[2\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{-6} \times 2= ?\]

OpenStudy (anonymous):

\[2\sqrt{-6}=2i \sqrt{6}\]

OpenStudy (anonymous):

yes because there is an invisble -1 which equals 1

OpenStudy (anonymous):

*i

OpenStudy (anonymous):

the \[\sqrt{-1}\] is an imaginary number and we call this \[i\]

OpenStudy (anonymous):

now multiply \[\sqrt{-6}\sqrt{-8}\]=?

OpenStudy (anonymous):

using the rule \[\sqrt{ab} =\sqrt{a}\sqrt{b}\]